OpenStudy (anonymous):

find the x-intercepts for the graphof the equation y=x^2+2x-3

OpenStudy (anonymous):

simply solve it by forming factors of x^2+2x-3 i.e., x^2+3x-x-3 x(x+3)-(x+3)=0 (x+3)(x-3)=0 x=-3,+3